OpenStudy (anonymous):

If a gas has a volume of 1 L at a pressure of 270 kPa, what volume will it have when the pressure is increased to 540 kPa? Assume the temperature and number of particles are constant. Show your work.

OpenStudy (anonymous):

P1V1=P2V2

OpenStudy (anonymous):

this is a variation of PV=nRT You are keeping everything on the right side constant, so you can ignore them

OpenStudy (anonymous):

really you only need to show that the pressure*volume at first is equal to the pressure*volume after

OpenStudy (anonymous):

Can you show me with the numbers

OpenStudy (anonymous):

(270)(1)=(540)(x) solve for x

OpenStudy (anonymous):

one thing to be careful of when doing these kinds of problems is be aware of your units

OpenStudy (anonymous):

how do you solve for x?

OpenStudy (anonymous):

divide both sides by 540

OpenStudy (anonymous):

0.5=1?

OpenStudy (anonymous):

.5=x not 1

OpenStudy (anonymous):

So, to show my work itll be: P1V1=P2V2 270(1)=540(x) 270=540 270 divided by 540 is .5 540 divided 540= 1 so .5 = x

OpenStudy (anonymous):

This step doesn't make sense 270(1)=540(x) 270=540 The x doesn't go away

OpenStudy (anonymous):

270=540(x) 540 divided by 270=.5 540 divided by 540=1 .5=x

OpenStudy (anonymous):

yes that works

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you're welcome =D